OpenStudy (anonymous):

Please help quick, it would save me big time! What is the slope, and what is the y-intercept of the line? y=5/4x+9 slope = ?/? y-intercept = ?

OpenStudy (texaschic101):

in y = mx + b form, the slope is in the m position and the y intercept is in the b position slope = 5/4 y int = 9

OpenStudy (anonymous):

Thanks! :D